1.) Time of flight (hint—you have to use vertical equations)

SHOW ALL WORK!

2.) Horizontal velocity

You shoot a nerf dart directly horizontal into the
Answer #1

1) Time of flight = T accleration due to gravity(a) = 9.8 initial velocity(u) = 0 downwards heigh(s) = 1m

So using s = ut + 0.5*a*t2 => 1 = 0 + 0.5*9.8*t2

=> t = sqrt(2*1/9.8) = 0.452 sec Answer

2) velocity = distance traveled/time_taken = 5.83 / 0.452 = 12.9 m/s Answer